The speed of a boat in still water is 15 km/h. The boat travels 40 km upstream and 40 km downstream in a total of 6 hours. What is the speed of the stream?
Answer by scott8148(6628)

time = distance / rate
total time is 6 hrs

let s="speed of stream"

[40 / (15 + s)] + [40 / (15 - s)] = 6

multiplying by (15 + s)(15 - s) ___ 600 - 40s + 600 + 40s = 1350 - 6s^2

0 = 150 - 6s^2 ___ 0 = 25 - s^2 ___ s = 5